# Advancements in Constitutive Model Calibration: Leveraging the Power of Full-Field DIC Measurements and In-Situ Load Path Selection for Reliable Parameter Inference

### v1

@@ 5 Discussion
In the construction of the ICC framework, certain assumptions and simplifications were made to ease the burden of algorithm development. Some of these constraints may be relaxed or adjusted in future applications; however, others are necessary in order to make ICC feasible with real-time data collection and model calibration, which is the future application space of ICC. These assumptions are bulleted and discussed below along with some suggestions of relaxing these constraints in the future where possible.
@@
    The posterior distribution was obtained with a Laplace approximation that assumes the posterior is Gaussian. Another method for approximating the posterior distribution which does not make such a strict assumption is Markov chain Monte Carlo (MCMC), which is widely used but very expensive. While the posterior is not guaranteed to be Gaussian, the Laplace approximation is obtained in less than a minute, while MCMC could easily take hours or more (conservatively) even with a fast surrogate model. Since the posterior is calculated after each load step, it is quickly realized that completing an entire experiment consisting of 5 or more load steps would not be feasible with such a computational demand. The Laplace approximation serves to quickly provide an estimate of parameter uncertainty which is then used in the EIG calculation to decide the next load step. It is possible that MCMC could result in a different adaptively selected load path if the posterior distribution is non-Gaussian; however, this possibility is secondary to efficient and practical decision-making. At the end of the data collection, MCMC could optionally be performed offline for a final calibration if desired.
@@
    The measurement noise variance for both the nodal displacements as well as the global load values was assumed to be known and i.i.d. In reality, noise from a load cell is a percent of the reading instead of constant. Additionally, while noise may be approximated to mimic what is seen in experiment, the statistical model would be more flexible if the noise variance were also inferred from the data. The incorporation of unknown and hetereoscedastic noise into the BOED framework was beyond the present scope, but may be included in future work.
@@
    A cruciform specimen in a planar biaxial load frame was selected for this synthetic experimental configuration. The corresponding load path tree had two branches per node, corresponding to tensile displacement in either the horizontal or vertical arms of the cruciform, with a fixed number of load steps and a fixed displacement increment for the load path tree. This was a simple starting point; however, ICC is not restricted to this tree structure. Other, more complex, and perhaps even more appropriate, branch options may include both positive and negative displacements, combinations of displacements applied in both directions simultaneously, a variable displacement increment, etc. Additionally, other deformation modes (i.e. tension-torsion) could be considered, or even a combination of planar biaxial and torsional loading given the capabilities of the load frame.
@@
    Quarter symmetry of the Cruciform Specimen
    Quarter symmetry was assumed in the circular region of interest of the cruciform to reduce the FE model size and conserve computational resources. While this is a reasonable assumption for the synthetic demonstration, a real-world lab environment may present more complex behavior.
@@
    Combining several QoIs for calibration is not foreign to engineering applications; both nodal displacements—transformed through PCA to yield corresponding singular values—and global load values were used to calibrate the material model in this work. However, a disparity in the number of measurements and magnitude of the values may lead to one QoI driving the likelihood and calibration disproportionately. While the number of displacement measurements ($>1,000\to 5$ after PCA) were on the same scale as the number of load measurements—1—for each directional component, there was a stark difference in the magnitude of the different QoIs—$1\times 10^{-6}-1\times 10^{-4}$ for the PCA singular values vs. $1\times 10^{3}$ N for the load. Working with kN instead of N would still present 4-6 orders of magnitude difference in the measurements. Thus, obtaining a proper balance between multiple QoIs in the likelihood is an important consideration. Table 5 shows the various contributions to the likelihood of the AAAAA load path at $\boldsymbol{\theta}^{\text{MAP}}$, revealing that the likelihood contribution from the PCA singular values is much greater than that from the load at the specified point. Future studies will further analyze the contribution of each QoI, and explore methodologies for ensuring a proper balancing of their contributions.
    $\boldsymbol{\theta}$ $\text{disp}^{\text{PCA}}_{X}$ $\text{disp}^{\text{PCA}}_{Y}$ $\text{load}_{X}$ $\text{load}_{Y}$ Log Likelihood
@@
    Table 5: Likelihood contributions from the multiple QoIs.
@@
    The load step selection was performed with a greedy (a.k.a myopic) decision-making protocol, where each load step was chosen based on the immediate expected gain in information—looking only one step ahead. The alternative is to make a globally optimal decision by calculating the EIG for the entire load path; however, this increases the cost of calculating the EIG exponentially. For example, at load step $t=2$, instead of calculating the EIG for 2 candidate load steps (Axis A vs. Axis B) in a mypocially-driven decision, the EIG would need to be calculated for the $n^{2}$ possible load paths which follow from the $n$ remaining number of load steps in a global decision. So, while the myopic algorithm is not guaranteed to result in the globally optimal load path, the resulting time savings make ICC possible in quasi-real-time and may yet result in the globally optimal decision (as in this example).
@@
    A fast Gaussian process (GP) surrogate replacement of the FE model was used for calibration and load step selection. A nice feature of GPs is that uncertainty surrounding the surrogate model can be incorporated into the inference problem. In this case, a simplified approach was taken, and the mean of the surrogate model was used without taking into account uncertainty of the surrogate model. The exclusion of the surrogate uncertainty was a simplification adopted for this initial demonstration, but future work may be expanded to include it. Additionally, because the surrogate model is trained on a grid of parameter values, the inference is restricted to this training space with truncated normal prior distributions. The parameter bounds used for training were chosen to be very conservative. However, if the support of the target posterior distribution falls outside the parameter bounds, it is likely the case that the chosen model-form is not appropriate for the specific application and should be revisited.
@@
In this phase of algorithm development, the ICC framework was constructed and implemented for the calibration of a single material model. However, the intent is to eventually use this framework for the calibration of multiple models as there is generally a level of uncertainty regarding the best model form for a given application. The Library of Advanced Materials for Engineering (LAMÉ) [34], part of the Sierra/SolidMechanics [35] finite element software, offers a modular framework for model development from which various model forms of differing degrees of fidelity may be explored. With the consideration of multiple models, there is the additional factor of determining how the step selection will proceed—this is the topic of future work.
@@
This work presented a demonstration of the ICC framework in a tightly-controlled synthetic environment. The synthetic demonstration was constructed to mimic the planned experiment as closely as possible, which is to deform an aluminum cruciform specimen in a planar biaxial load frame and perform calibration with full-field DIC data. A first demonstration of the ICC framework has been performed experimentally, and a primary challenge moving to the real-world laboratory conditions is machine control. The surrogate models were constructed with some assumed boundary conditions, and reliable use of these models requires that the realized experimental boundary conditions match the ones used for surrogate training as closely as possible. With some adjustments to the ICC framework [49], the strict boundary conditions were closely met and ICC was completed successfully. However, future demonstrations may require that surrogates be trained over a range of boundary conditions in order to relax the requirements on machine control. Additionally, in this synthetic demonstration, the calibration and load step selection was performed in the absence of model-form error since the same model was used to both generate the data and in calibration. In a real-world application, no model can perfectly capture the the true behavior of the material. If there is an observed systematic discrepancy between the data and the model, model-form error will need to be taken into account in order to obtain credible results in the calibration.
@@ 6 Conclusions
Computational simulation is relied upon for high-consequence engineering decisions; in solid mechanics simulations, a foundational element is a credible constitutive (material) model which can compute the stresses in a material under a specified loading condition. The usefulness of these simulations for decision-making not only relies on the credibility of the calibrated material-dependent parameters—and inherently the appropriateness of the data collected through material characterization—but also on the ability to efficiently achieve the credible calibration for the material model. Both of these considerations support the credible use of computation simulations at early stages in design cycles.
Historically, the material characterization and model calibration workflow has been sequential, inflexible, and potentially incomplete. First, relevant phenomenological features are assumed a priori based on subject matter expertise. Next, a suite of characterization experiments are designed and performed based on those assumptions. Most often, the experiments target specific stress states (i.e. tension, shear, etc.) and collect only global force-displacement data for calibration. Then, typically the material model is calibrated using a deterministic process that provides a best-fit parameter estimate with no uncertainty metrics. Finally, a distinct validation effort may—or may not—be performed. Only at the end of this process is the accuracy of the calibration ascertained. If the calibration is inadequate, then the entire process must be repeated, and it can take months or even years before a satisfactory calibration is obtained.
@@
This work brings together several recent advancements made in material characterization and model calibration into an improved workflow—Interlaced Characterization and Calibration (ICC)—that improves not only the process of obtaining a reliable material model calibration, but also the quality of the calibration itself. The ICC framework achieves these improvements by 1) replacing uniaxial test specimens with a cruciform specimen that reveals complex stress states, 2) replacing global load-displacement curves with full-field displacement data, 3) replacing deterministic calibrations and best-fit parameter estimates with Bayesian inference, which produces a distribution quantifying parameter uncertainty, 4) adaptively selecting the load steps—within the confines of a load path tree—as the cruciform is deformed through Bayesian optimal experimental design (BOED) instead of relying on subject matter expertise to select a load path a priori, and 5) performing steps 1-4 in a quasi-real-time feedback loop. The result is that informative experimental data is collected in a highly-efficient manner to yield a credible material model that can be used for early decision-making.
@@
A synthetic demonstration of the ICC framework was presented which considered the deformation of cruciform specimen in a biaxial load frame to calibrate an elastoplastic constitutive model. In the exemplar problem, calibration was performed with a surrogate replacement of the expensive FE model, and full-field displacement data and global load data were synthetically generated for calibration. The displacement field was dimensionally reduced via principal component analysis (PCA) to improve numerical stability and aid in the construction of surrogate models, and inference was performed with the resulting PCA singular values. A non-alternating AAAAA load path (pulling on the same axis for all five load steps) was preferred by the algorithm and resulted in a calibration that had the least parameter uncertainty compared to all other load paths—a non-intuitive result. This outcome reinforced the utility of BOED algorithms to identify load paths that collect more informative data than load paths selected a priori by subject matter experts.
@@
The tools developed for this synthetic demonstration make a significant step in moving towards a real-time demonstration of the ICC framework in the lab. By using this approach, high-fidelity material models can be calibrated with a great increase in efficiency over current methods. Additionally, the proposed approach improves the quality of the calibrations with guided data collection and quantified uncertainty. These results improve the state-of-the-art in standard calibration procedures.

### v3

@@ 5 Discussion
In the construction of the ICC framework, certain assumptions and simplifications were made to ease the burden of algorithm development. Some of these constraints may be relaxed or adjusted in future applications; however, others are necessary in order to make ICC feasible with real-time data collection and model calibration, which is the future application space of the ICC framework. These assumptions are bulleted and discussed below along with some suggestions of relaxing these constraints in the future where possible.
@@
    The posterior distribution was obtained with a Laplace approximation that assumes the posterior is Gaussian. Another method for approximating the posterior distribution which does not make such a strict assumption is Markov chain Monte Carlo (MCMC), which is widely used but very expensive. While the posterior is not guaranteed to be Gaussian, the Laplace approximation is obtained in less than a minute, while MCMC could easily take hours or more, conservatively, even with a fast surrogate model. Since the posterior is calculated after each load step, it is quickly realized that completing an entire experiment consisting of 5 or more load steps would not be feasible with such a computational demand. The Laplace approximation serves to quickly provide an estimate of parameter uncertainty which is then used in the EIG calculation to decide the next load step. It is possible that MCMC could result in a different adaptively selected load path if the posterior distribution is non-Gaussian; however, this possibility is secondary to efficient and practical decision-making. At the end of the data collection, MCMC could optionally be performed offline for a final calibration if desired. Other methods with reduced time requirements for posterior approximation compared to MCMC include variational inference [71, 72] and sequential Bayesian methods [73, 74], which may be considered in the future.
    In order to support the use of the Laplace approximation in this work, a validation of the posterior approximation was obtained by performing an MCMC simulation with the collected data, and the results are included in B. Results showed that the posterior distribution was well-approximated with a Gaussian distribution in this case. The posterior obtained via MCMC took approximately 1.5 hours to obtain; however, this time requirement could increase significantly in future applications if the posterior is difficult to sample from. In comparison, posterior approximation took less than a minute with the Laplace approximation.
@@
    The measurement noise variance for both the nodal displacements as well as the global load values was assumed to be known and i.i.d. In reality, noise from a load cell is a percent of the reading instead of constant, and displacement field data has a noise structure that is spatially correlated [31]. The statistical model would be more flexible and provide more reliable uncertainty metrics in the experimental application if the noise variance were also inferred from the data. The incorporation of unknown hetereoscedastic noise into the BOED framework has been left to future work, and the correlated noise of the field data can be addressed by performing inference in the spectral domain, where PCA singular values offer a decorrelated representation as addressed in the Inference in Field vs. PCA Space discussion point.
@@
    A cruciform specimen in a planar biaxial load frame was selected for this synthetic experimental configuration. The corresponding load path tree had two branches per node, corresponding to tensile displacement in either the horizontal or vertical arms of the cruciform, with a fixed number of load steps and a fixed displacement increment for the load path tree. This was a simple starting point; however, ICC is not restricted to this tree structure. Other, more complex, and perhaps even more appropriate, branch options may include both positive and negative displacements, combinations of displacements applied in both directions simultaneously, a variable displacement increment, etc. Additionally, other deformation modes—such as tension-torsion, or even a combination of planar biaxial and torsional loading given the capabilities of the load frame—could be considered.
@@
    Symmetry of the Cruciform Specimen
    Eighth-symmetry was used in constructing the FE model to reduce the size and computational cost. While this is a reasonable idealization for the synthetic demonstration, laboratory results often deviate from such approximations. For instance, recent investigations on complex specimens [35] have demonstrated that utilization of ideal in place of actual boundary conditions can impact predicted results. Therefore, when used in practice, such symmetries may not be exploitable.
@@
    In engineering applications, it is common to combine multiple QoIs for calibration to capture various aspects of material behavior. For example, this work utilized both full-field displacements and global load values to calibrate a material model. While the calibration could have been performed using only the load data or only the displacement data, using both data sources and combining the unique information they contain greatly improves the calibration. Table 5 shows the posterior summaries using solely load data, solely displacement or both sources combined. The expected values indicate that all options effectively infer parameter values from the available data. However, when both sources are used together, parameter uncertainty is reduced. Figure 18 visually compares the different cases, clearly demonstrating the advantages of incorporating both data sources in the calibration process.
    The posterior uncertainty significantly decreases when adding displacement data to the load-only calibration (a). Conversely, transitioning from the displacement-only calibration to the combined calibration with load data (b) results in a much smaller reduction in posterior uncertainty. This difference highlights the varying information content of each QoI: the displacement data which has over 1,000 measurements provides rich information, while the load data which provides a single measurement contributes less. While balancing contributions from multiple QoIs may be necessary in some cases—for example, to avoid over-fitting—both data sources in this example, as shown in Fig. 18 and supported by posterior predictive checks in Sec. 4.3, meaningfully contributed to the inference.
    Load path Data Space $\boldsymbol{\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}}$ $\boldsymbol{\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}}$ Gen. Var.
    $\boldsymbol{(1\times 10^{-4})}$ $\boldsymbol{(1\times 10^{-14})}$
@@
    Disp. [42.49, 13.64, 14.28, 11.17] [ 9.4, 3.2, 19.0, 6.9 ] 2.06
    Load [42.53 13.70, 14.18, 11.23] [ 30.5, 275.8, 1922.4, 32.4 ] $3.06\times 10^{4}$
    $\boldsymbol{\theta}^{true}=[42.51,13.63,14.35,11.19]$
    Table 5: Posterior summaries obtained for load path AAAAA using solely load data, solely displacement data, or both data sources combined.
[…]
    [figure]style=plain,subcapbesideposition=top,font=small
[…]
    Figure 18: Posterior contour comparisons for a) calibrating with both full-field data and load data vs. solely load data, and b) calibrating with both full-field data and load data vs. solely displacement data for load path AAAAA. Marginal posterior densities are plotted on the diagonals and joint-marginal densities on the off-diagonals. Marginal densities for the calibration performed with both data sources are plotted with a green line and with a single data source with an orange line. $\boldsymbol{\theta}^{true}$ is indicated with a dashed pink line and star. The contours shown for the combined calibration are the same as those shown in Fig. 10.
@@
    The load step selection was performed with a greedy (a.k.a. myopic) decision-making protocol, where each load step was chosen based on the immediate expected gain in information—looking only one step ahead. The alternative is to make a global decision by calculating the EIG for the entire load path; however, this increases the cost of calculating the EIG exponentially. For example, at load step $t$, instead of calculating the EIG for 2 candidate load steps in a mypocially-driven decision (load step A vs. load step B), the EIG would need to be calculated for the $(\mathcal{T}-t)^{2}$ possible load paths which follow from all possible future load steps in a global decision. So, while the myopic algorithm is not guaranteed to result in the globally preferred load path, the resulting time savings make ICC possible in quasi-real-time and may yet result in the globally preferred decision, as in this example. As previously mentioned, sequential OED approaches [70] may allow for the selection of the globally preferred full load path with a manageable computational demand. However, due to the non-Markovian nature of materials, such an approach would be challenging in this application and exploration of its use is reserved for future endeavors.
@@
    A fast Gaussian process (GP) surrogate replacement of the FE model was used for calibration and load step selection. A nice feature of GPs is that uncertainty surrounding the surrogate model can be incorporated into the inference problem. In this case, a simplified approach was taken, and the mean of the surrogate model was used without taking into account uncertainty of the surrogate model, but future work may be expanded to include the uncertainty. In general, the mean absolute error between the surrogate output and the model (evaluated at locations in the parameter space reserved for testing) was less than 0.003% of the load and the less than 0.1% of the displacements (which included any error from the PCA transformation as well). Thus, the error due to the surrogate model approximation was significantly less than the expected measurement noise for both QoIs. Additionally, because the surrogate model is trained on a grid of parameter values, the inference is restricted to this training space with truncated normal prior distributions. The parameter bounds used for training were chosen to be very conservative. However, if the support of the target posterior distribution falls outside the parameter bounds, it is likely the case that the chosen model-form is not appropriate for the specific application and should be revisited.
  • 8.
    Inference in Field vs. PCA Space
    When calibrating a model using full-field data, the choice between performing inference in PCA space versus the original field space presents both advantages and disadvantages. Calibrating in field space allows for the direct incorporation of experimental data, enhancing the accuracy of parameter estimates. The simplified model used in this work assumed the noise was i.i.d. While this is an appropriate assumption for synthetically generated data, when moving to a live demonstration of ICC with experimental data, correlated or structured noise in the field data [31] may lead to underestimated uncertainty or biased parameter estimates. In such a case, a more complex model which accounts for any such noise characteristics may be necessary. Conversely, PCA space offers the benefits of dimension reduction, decorrelation of the noise structure and simplification of the calibration process, significantly speeding up tasks like the EIG approximation. For instance, in this synthetic example, calculating the EIG with field data took approximately 40 minutes, while using PCA singular values reduced this time to just 30 seconds. However, since the PCA modes are derived from model evaluations devoid of noise, they may not fully capture the noise structure of the experimental data, potentially contributing to model form error and parameter bias. To fully leverage the benefits of PCA-based inference while mitigating potential bias, it may be necessary to incorporate the noise structure of the experimental data directly into the PCA basis or to model the residual noise explicitly in the spectral domain.
    Calibrations in both the field space and the PCA space were performed in this work, with comparisons of the posterior summaries and contours shown in Table 6 and Figure 19. The results are very similar for this example; however, the posterior contours and marginal variances indicate that a higher posterior probability is placed on the true values with the field data, resulting in a less-biased parameter expected value. As a result, the main body of the work discusses inference in the field space, allowing a direct incorporation of experimental data and reducing parameter bias.
    While the time to calculate the EIG in the field space (40 min at each step) was not prohibitive for this synthetic demonstration, it would be during a live demonstration. A comparison of computing the EIG in the PCA space (while maintaining the calibration in the field space) vs. the field space in Table 7 reveals that the two methods yield identical approximations—to the 11th digit. Thus, when moving to an experimental demonstration, the preferred method is to perform the EIG approximation in the PCA space. Future decisions of performing inference in the field or PCA domain must reflect a careful consideration of the trade-offs between accuracy, computational efficiency, and the need to capture the inherent complexities of the experimental data.
    Load path Data Space $\boldsymbol{\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}}$ $\boldsymbol{\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}}$ Gen. Var.
    $\boldsymbol{(1\times 10^{-4})}$ $\boldsymbol{(1\times 10^{-15})}$
[…]
    PCA [42.50, 13.67, 14.28, 11.21] [ 1.6, 2.5, 15.9, 5.3 ] 2.37
    $\boldsymbol{\theta}^{true}=[42.51,13.63,14.35,11.19]$
    Table 6: Posterior summaries obtained for load path AAAAA using the load plus displacement data in the field space vs. the reduced PCA space (i.e. PCA singular values).
    Figure 19: The posterior contours for load path AAAAA using full-field displacement data and PCA singular values. Marginal posterior densities are plotted on the diagonals and joint-marginal densities on the off-diagonals. Marginal densities for the full-field calibration are plotted with a green line and for the PCA calibration with an orange line. $\boldsymbol{\theta}^{true}$ is indicated with a dashed pink line and star. The contours shown for the calibration in the field space are the same as those shown in Fig. 10.
    Load path Load Step Field EIG PCA EIG
    [A, B] [A, B]
[…]
    2 [8.578, 4.144] [8.578, 4.144]
    3 [4.388, 2.643] [4.388, 2.643]
    4 [2.423, 1.344] [2.423, 1.344]
    5 [2.139, 1.125] [2.139, 1.125]
    Table 7: A comparison of the EIG calculated in the field space vs. the PCA space for both load step options (A and B) and for each load step except the first; the first load step was pre-determined to be A. The values are identical to the 11th digit, with the PCA-based calculations taking only a fraction of the time as the field-based calculations.
@@
In this phase of algorithm development, the ICC framework was constructed and implemented for the calibration of a single material model. However, the intent is to eventually use this framework for the calibration of multiple models as there is generally a level of uncertainty regarding the best model form for a given application. The Library of Advanced Materials for Engineering (LAMÉ) [37], part of the Sierra/SolidMechanics [38] finite element software, offers a modular framework for model development from which various model forms of differing degrees of fidelity may be explored. With the consideration of multiple models, there is the additional factor of determining how the step selection will proceed—this is a topic of future work.
Preferred Load Path Insights
In this study, the AAAAA load path was identified as the preferred load path for generating the most informative data for calibration, leading to a greater degree of confidence in the parameter estimates. The preference of the AAAAA load path over all others is influenced by a number of factors, including the specimen geometry, model form, relative impact of material phenomenologies and the structure of the load path tree. Even with the same experimental set-up, the resultant preferred path may change depending on the material response. For instance, the preference of the AAAAA load path for a cruciform specimen and the calibration of an isotropic model contrasts with the results presented in Ricciardi et al. [30], which indicated that an alternating ABABA load path was more advantageous for a material point undergoing similar deformation and the calibration of an anisotropic model, assuming the same load path tree in both examples. This discrepancy underscores the complexity of the issue and suggests that while optimized experimental strategies implicitly assume that richer information content leads to better calibration, understanding the reason behind the superiority of one load path over another is non-trivial. Such activities, while important, are left to future studies.
@@
This work presented a demonstration of the ICC framework in a tightly-controlled synthetic environment. The synthetic demonstration was constructed to mimic a planned experiment as closely as possible, which is to deform an aluminum cruciform specimen in a planar biaxial load frame and perform calibration with full-field DIC data. A first demonstration of the ICC framework has been performed experimentally, and a primary challenge moving to the real-world laboratory conditions was found to be machine control. The surrogate models were constructed under assumed boundary conditions, and reliable use of these models requires that the realized experimental boundary conditions match the ones used for surrogate training as closely as possible. With some adjustments to the ICC framework [52], the strict boundary conditions were closely met and ICC was completed successfully. However, future demonstrations may require that surrogates be trained over a range of boundary conditions in order to relax the requirements on machine control. Additionally, in this synthetic demonstration, the calibration and load step selection were performed in the absence of model-form error since the same model was used to both generate the data and perform the calibration. In a real-world application, no model can perfectly capture the the true behavior of the material. If there is an observed systematic discrepancy between the data and the model, model-form error will need to be taken into account in order to obtain credible uncertainty estimates from the calibration.
Calibration Efficiency
In this synthetic demonstration, each load step required approximately 40 minutes in total, with less than a minute needed to approximate the posterior and around 40 minutes to estimate the EIG for both load step options. Thus, the entire synthetic demonstration of the ICC framework took less than 3.5 hours for 5 total load steps. Transitioning towards implementing the ICC framework with real-time data collection and model calibration, it is anticipated that material characterization and model calibration can be completed within a single day. Considering the setup time for both the experiments and the surrogates, it is projected that a fully developed ICC framework could accomplish the entire characterization and calibration process (excluding the time required for specimen machining) in just 1 to 2 weeks. In contrast, the traditional approach typically allocates one week for each experimental setup, which could extend the timeline to 6 to 7 weeks for a complete test matrix. Additionally, an extra 2 to 3 weeks is often needed for data reduction, FEA meshing and setup and calibration, resulting in an overall duration of 8 to 10 weeks. Therefore, the ICC paradigm has the potential to achieve a remarkable 5 to 10 times increase in efficiency compared to traditional methodologies.
@@ 6 Conclusions
This work brings together several recent advancements made in material characterization and model calibration into an improved workflow—Interlaced Characterization and Calibration (ICC)—that improves not only the process of obtaining a reliable material model calibration, but also the quality of the calibration itself. The ICC framework achieves these improvements by 1) replacing uniaxial test specimens with a cruciform specimen that reveals complex stress states, 2) replacing global load-displacement curves with full-field displacement data, 3) replacing deterministic calibrations and best-fit parameter estimates with Bayesian inference, which produces a distribution quantifying parameter uncertainty, 4) adaptively selecting the load steps—within the confines of a load path tree—as the cruciform is deformed through Bayesian optimal experimental design (BOED) and 5) performing steps 1-4 in a quasi-real-time feedback loop. The result is that informative experimental data is collected in a highly-efficient manner to yield a credible material model that can be used for early decision-making.
@@
A synthetic demonstration of the ICC framework was presented which considered the deformation of a cruciform specimen in a biaxial load frame to calibrate an elastoplastic constitutive model. In the exemplar problem, calibration was performed with a surrogate replacement of the expensive FE model, and full-field displacement data and global load data were synthetically generated for calibration. The displacement field was dimensionally reduced via principal component analysis (PCA) to aid in the construction of surrogate models; the surrogate predictions of the PCA singular values were then used to reconstruct the displacement field to perform inference and calculated the EIG in the field space. A non-alternating AAAAA load path (pulling on the same axis of the cruciform for all five load steps) was preferred by the algorithm and resulted in a calibration that had the least parameter uncertainty compared to all other load paths—a non-intuitive result. This outcome reinforced the utility of BOED algorithms to identify load paths that collect more informative data than load paths selected a priori by subject matter experts.
@@
The tools developed for this synthetic demonstration make a significant step in moving towards a demonstration of the ICC framework in the lab with real-time data collection. By using this approach, high-fidelity material models can be calibrated with a great increase in efficiency over traditional methods. Additionally, the proposed approach improves the quality of the calibrations with guided data collection and quantified uncertainty. These results improve the state-of-the-art in standard calibration procedures.